OpenStudy (anonymous):

log(2)(x+4)+log(2)(x-2)=4 x=___? I got 1 as my answer, right??

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

lol, i suck at this....still

myininaya (myininaya):

satellite i need your help

OpenStudy (anonymous):

x cannot be 1 because you have \[\log_2(x-2)\]

myininaya (myininaya):

when you are done

OpenStudy (anonymous):

and 1-2=-1 and you cannot take the log of a negative number

OpenStudy (anonymous):

ok hold on lets finish this one

OpenStudy (anonymous):

you get \[log_2((x+4)(x-2))=4\] \[(x+4)(x-2)=2^4=16\]

OpenStudy (anonymous):

16???? I was waaaaay off!!!!

OpenStudy (anonymous):

not done yet!

OpenStudy (anonymous):

ok, I need to find x.. ok

OpenStudy (anonymous):

\[x^2+2x-8=16\] \[x^2-2x-24=0\] \[(x+6)(x-4)=0\] \[x=4\]

OpenStudy (anonymous):

ignore -6

OpenStudy (anonymous):

thnx, u rock

OpenStudy (anonymous):

yw